UT 20130916 Technical Observing - AGw2 Calibration
Observer: JHill, DMiller (Tucson)
Telescope Operator: SAllanson (LBTO)
Software Support: DCox (Tucson)
AO Support: DMiller (Tucson)
Telescope Support: (LBTO)
Instrument Support: JPower (LBTO)
Testing DX Adaptive Optics for 2 hours at the beginning - the AO system is working nominally. Then we closed for weather.
01:40 I'm completely shocked that we've opened the dome at the normal time. However we are still dancing with the humidity limit.
DX Adaptive Optics Testing
02:00 Arcetri guys are testing the new IM on DX M2 @ RFBG.
02:40 Simone reports that things are working well with 60- 90% Strehl at H-band - apparently limited by vibrations. There were "no surprises" with
the inner edge of the shell near the mousehole - forces are within range.
03:40 Close for a threatening cloud off to the east, and then the humidity rises.
Today we progressed in the BIN1 calibrations measuring two
different type of reconstructors associated with two different modal basis.
We were ready to test them on sky and J. Hill let us go
from 7:40 to about 9:20 when the telescope was closed for high humidity.
The good news is that the reconstructor we tested (derived from a modal
base with the first ring of actuators slaved) achieved SR between 60 and
87% SR in H band showing no particular sign of excess force on the actuator
of the internal ring. So it seems that the modal basis and optical masking
and WFS digital masking we introduced are able to deal with the warping of
TS4 in the central region. We wanted to test a second reconstructor but
observation stopped for high humidity. Considering the good result on
sky, we will move to measure calibration for the BIN2 mode.
Had we opened, we would have applied to DX M2 hexapod about -0.8 of the repointing that Doug measured. ---> Apply Tip=-25 Tilt=-100 arcsec
05:45 Thunderstorm forms over basecamp.....
We are occupying our time testing pseudomonocular mode observations with closed-dome.
- 15 Sep 2013